An architect drew the blueprint for a new office building. He used a scale in which 1 inch represents 7.5 feet. The floor of an office in the building will have actual dimensions of 20 feet by 20 feet. What will be the dimensions on the blueprint?
Answer by TimothyLamb(4379)

i/f = 1/7.5
i/20 = 1/7.5
i = 20*1/7.5
i = 2.6666666666
---
answer:
2.67 inches by 2.67 inches
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---